The voltage across in an air-filled parallel-plate capacitor is measured to be 85.0V.
When a dielectric is inserted and completely fills the space between the plates, the voltage drops to 25.0V
(i) What is the dielectric constant of the inserted material?
1
Expert's answer
2021-02-18T18:56:37-0500

(i) The initial capacitance of an air-filled parallel-plate capacitor can be written as follows:


"C_i=\dfrac{\epsilon_0A}{d}."

The final capacitance of the parallel-plate capacitor filled with dielectric can be written as follows:


"C_f=\dfrac{\kappa\epsilon_0A}{d}."

Let's divide "C_f" by "C_i":


"\dfrac{C_f}{C_i}=\kappa."

From the other hand,


"C=\dfrac{Q}{\Delta V}."

Therefore, we can write:


"\kappa=\dfrac{C_f}{C_i}=\dfrac{\dfrac{Q}{\Delta V_f}}{\dfrac{Q}{\Delta V_i}}=\dfrac{\Delta V_i}{\Delta V_f}=\dfrac{85\ V}{25\ V}=3.4."